OpenStudy (anonymous):

A 400 g soccer ball lying on ground is kicked and flies off at 25 m/s. If the duration of the impact was 0.010 s, what was the average force on the ball?

OpenStudy (anonymous):

average force = Change in momentum / time of impact

OpenStudy (anonymous):

average force= 25 / 0.010 is this correct?

OpenStudy (aaronq):

you're missing the mass in the equation, momentum=mass*velocity

OpenStudy (anonymous):

average force = (25)(400) / 0.010 = 1000000

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

think properly danielle.. that is still not right.. what did you miss? (check units)

OpenStudy (anonymous):

do i need to convert 400 g to kg?

OpenStudy (anonymous):

That is S.I units right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so then you have to :)!

OpenStudy (anonymous):

ok. thanks!

OpenStudy (anonymous):

no problem :)